If you wanted to fill a rubber balloon to an internal volume of 3.750 L and a pressure of 128.0 kPa in a 24.60°C room with hydrogen gas from the reaction of metal and acid ( Be + HIO3), what is the minimum quantity (in mL) of a 6.500M solution of your acid you would need to generate enough hydrogen gas? (Assume you have more than enough metal)


1
Expert's answer
2021-06-23T05:16:59-0400

PV = nRT

P = 128×0.00985 = 1.261 atm

T = 24.6+273 = 297.6K

n = 6.5

R = 8.314

V = 6.5 × 8.314 × 297.6/1.261

= 12753.85ml - 3750

= 9003.85ml